OpenStudy (bloomlocke367):

5 years ago Kate was 5 times as old as her Son. 5 years hence her age will be 8 less than three times the corresponding age of her Son. Find their ages.

OpenStudy (bloomlocke367):

@bohotness

OpenStudy (bloomlocke367):

@dan815 @.Sam. @perl

OpenStudy (kropot72):

The following equations can be written, where k is the present age of Kate and s is the present age of her son: k - 5 = 5(s - 5) ............(1) k + 5 = 3(s + 5) - 8 .....(2) Now you need to solve these simultaneous equations.